# From Intentions to Techniques: A Comprehensive Taxonomy and Challenges in Text Watermarking for Large Language Models

### v1

@@ 2.2.3 Ad-Hoc Addition
Rule-based substitutions and watermark additions at the embedding level are the most popular ways to add watermarks. However, multiple addition techniques do not fit into any of the two categories. We bucket these methods into Ad-Hoc addition methods and list a few methods that we found relevant.
First, Por et al. (2012); Sato et al. (2023) insert Unicode space characters in various text spacings. For example, Sato et al. (2023) proposes three different methods: WhiteMark, VariantMark, and PrintMark. WhiteMark operates by substituting whitespace characters with alternate Unicode whitespace characters, such as replacing U+0020 with U+2004. Variantmark emerges as a specialized watermarking technique tailored for Chinese, Japanese, and Korean texts. Leveraging Unicode’s variation selectors, Variantmark embeds secret messages by replacing Chinese characters with their variants. Printmark addresses the challenge of watermarking printed texts through nuanced strategies that subtly alter text appearance. It employs ligatures, varying whitespace lengths, and utilizing variant characters.
Another work by Atallah et al. (2001) introduces three unique syntax transformations for message encoding— Adjunct Movement, Clefting, and Passivization. For instance, Adjunct Movement involves relocating adjuncts within a sentence, as demonstrated by the variability in positioning the word ’quickly’ in "She quickly finished her homework." Clefting highlights a specific sentence part, typically the subject, such as transforming "The chef cooked a delicious meal" into "It was the chef who cooked a delicious meal" to emphasize ’the chef.’ Passivization, on the other hand, changes active sentences with transitive verbs into passive voice, like transforming "The teacher graded the exams" into "The exams were graded by the teacher." Each transformation corresponds to a unique message bit: Adjunct Movement to 0, Clefting to 1, and Passivization to 2.
Lastly, Sun et al. (2023) involves changes in the operators of the code based on adaptive semantic-preserving transformations.
@@
Text Completion Colossal Clean Crawled Corpus (C4) Raffel et al. (2020), Dbpedia Class Auer et al. (2007), WikiText-2 Merity et al. (2016) Kirchenbauer et al. (2023), Kuditipudi et al. (2023), Liu et al. (2023a), Munyer and Zhong (2023), Yoo et al. (2023b), Liu et al. (2023b), Fairoze et al. (2023), Ren et al. (2023), Hou et al. (2023), Qu et al. (2024)
Post-watermark text similarity analysis WikiText-2, Workshop on Statistical Machine Translation (WMT14) Bojar et al. (2014), Internet Movie Database (IMDb) Maas et al. (2011), AgNews Zhang et al. (2015), Dracula, Pride and Prejudice, Wuthering Heights Gerlach and Font-Clos (2020), CNN/Daily Mail Nallapati et al. (2016), Human ChatGPT Comparison Corpus (HC3) Guo et al. (2023), C4, Reuters Corpus Lewis et al. (2004), ChatGPT Abstract Nicolai Thorer Sivesind (2023), Human Abstract Nicolai Thorer Sivesind (2023) Yang et al. (2022), He et al. (2022a), He et al. (2022b), Yoo et al. (2023a), Yang et al. (2023b), Sato et al. (2023), Topkara et al. (2006a), Zhang et al. (2023)
@@
Code Generation CodeSearchNet (CSN) Husain et al. (2019), HUMANEVAL Chen et al. (2021), Mostly Basic Python Programming (MBPP), MBXP Athiwaratkun et al. (2023), DS-1000 Lai et al. (2023) Lee et al. (2023), Li et al. (2023), Yang et al. (2023a)
Question Answering OpenGen Krishna et al. (2024), Long Form Question Answering (LFQA) Krishna et al. (2024) Zhao et al. (2023a), Yoo et al. (2023b), Qu et al. (2024)
@@ 2.3 Evaluation
A wide variety of datasets have been used to evaluate the performance of watermarking approaches, limiting our ability to extract generalized conclusions about their performance. Different benchmarks focus on selected downstream tasks to validate watermarking capabilities, and we provide a detailed breakdown of the datasets utilized in Table 3. We observe that there are a large number of evaluation datasets focusing on text completion and post-watermarking text similarity tasks. The downstream task descriptions are provided below.
@@
Text Completion Task: This task involves giving the LLM a portion of text from the dataset as a prompt and then asking it to complete the text. The generated completion is then compared with the human completion or the portion of the dataset not provided as the prompt.
Post-watermark text similarity analysis: In this task, given an initial text $X$, watermarking is applied to $X$ to produce a modified text $X^{\prime}$. An example could be a rule-based substitution with synonyms or spelling replacements. The comparison is then made between $X$ and $X^{\prime}$, with $X$ and $X^{\prime}$ on the basis of distinctions in length, semantic, and other linguistic features.
Other Downstream Tasks: For these tasks, given the same initial prompt $X$, the LLM’s generated response $Y$ (before watermarking) is compared with the response $Y^{\prime}$ (after watermarking). This evaluates how watermarking affects the LLM’s output.
@@ 2.4 Adversarial attacks on watermarking techniques
Figure 5: An example of an adversary performing a de-watermarking attack on a green-red list-based watermarking technique. The original partitioning contains a higher proportion of green tokens as compared to the partitioning after adversarial paraphrasing.
@@
    Text insertion attacks involve adding additional tokens or text segments to the original output of a watermarked LLM generation. For example, on watermarking methods with green-red lists Kirchenbauer et al. (2023); Zhao et al. (2023b); Takezawa et al. (2023), an attacker could add additional tokens from the red list leading to the obfuscation of the watermarking method.
@@
    Text deletion attacks involve the removal of tokens or text segments from the original watermarked output of an LLM and modifying the rest of the tokens to fit the output. Coming back to the example of green-red list methodologies, this means removing some of the green list tokens from the output and modifying the red list tokens in the output. These techniques often require knowledge of the vocabularies belonging in each of the two lists in green-red lists.
@@ 3 Discussion and Open Challenge
We describe the open challenges to watermarking and outline "good to have" criteria while developing new techniques to protect intellectual property ownership. They are as follows:
@@
One of the critical challenges in the field is the lack of comprehensive evaluation against a diverse range of de-watermarking attacks. While many researchers focus on developing robust watermarking techniques, there is often insufficient emphasis on systematically red-teaming these methods against multiple attacking scenarios.
@@
Watermarks modify the model output distributions; techniques that are robust to de-watermarking often have greater variations in watermarked outputs compared to clean outputs leading to a potential trade-off between de-watermarking and LLM factuality. Despite this potential trade-off, there is a lack of analysis on how watermarking techniques affect the output inaccuracies or hallucinations. After training or fine-tuning LLMs with specific watermarking techniques, there is often insufficient examination of whether these methods introduce or exacerbate inaccuracies. We advocate for factuality evaluations post-watermarking.
Compatibility to various NLP downstream tasks
Important task types like Story Generation, Text Classification etc. are under-explored.
Enhanced interpretablity
@@
We urge the community to work on human perception of LLMs when interacting with different safety principles. User perception of LLMs may change with differences in output distributions. Furthermore, safety practices may enable AI acceptance and adoption among the masses.
@@ 4 Conclusion
In this paper, we analyze representative literature in the field and provide a comprehensive taxonomy for digital watermarking techniques for both LLM-generated and human-written text. The taxonomy categorizes watermarking techniques using four primary categories, namely - intention of the method, data used for evaluation, watermark addition, and removal.
Our work not only identifies and clusters existing watermarking methods but also brings to light key open challenges and research gaps in the field. These challenges include the need for more rigorous testing against diverse de-watermarking attacks, the establishment of standardized benchmarks for fair and consistent comparison of different techniques, and a deeper understanding of how watermarking impacts the factuality and accuracy of LLM outputs. Furthermore, we emphasize on the importance of developing watermarking techniques that are resilient to adversarial attacks, enhance interpretability, and maintain compatibility across various NLP downstream tasks.
We envision this research to serve as a reference for policymakers, safety practitioners, and end users; facilitating the adoption of robust digital watermarking practices and promoting responsible AI use.
@@ 5 Limitations
Limitations to our work are as follows: (1) We do not include detailed insights into metrics for success rate (accuracy of detecting watermarked texts), text quality (perplexity and semantics), NLP task-specific evaluation, and robustness (detectability of watermarks after removal attacks). (2) We don’t demonstrate the mathematical analysis of different watermarking techniques. (3) We do not cover all different task deployment scenarios for the watermarking techniques discussed.

### v2

@@ 2.2.3 Ad-Hoc Addition
Unlike popular watermarking methods like rule-based substitutions, which have strict, global definitions for modifications within a sentence (like synonym/spelling replacement and triggers), the ad-hoc addition methods use task-specific local guidelines for changes to the sentence structure. We bucket these methods into Ad-Hoc addition methods and list a few relevant methods.
First, Por et al. (2012); Sato et al. (2023) embed watermarks by inserting Unicode spaces in the text. Sato et al. (2023) introduce three methods: WhiteMark replaces whitespace with alternate Unicode spaces (e.g., U+0020 to U+2004), VariantMark uses Unicode variation selectors to embed messages in Chinese, Japanese, and Korean texts by substituting characters with variants, and PrintMark alters text appearance for printed media through ligatures, varied spaces, and character variants. Another work introduces three unique syntax transformations for message encoding— Adjunct Movement, Clefting, and Passivization Atallah et al. (2001). For instance, Adjunct Movement involves relocating adjuncts within a sentence, as demonstrated by the variability in positioning the word ’quickly’ in "She quickly finished her homework." Clefting highlights a specific clause, typically the subject, such as transforming "The chef cooked a delicious meal" into "It was the chef who cooked a delicious meal" to emphasize ’the chef.’ Passivization, however, changes active sentences with transitive verbs into passive voice, transforming "The teacher graded the exams" into "The exams were graded by the teacher." Sun et al. (2023) apply semantic-preserving code transformations by modifying operators.
Overlapping categories
Some papers span multiple categories within the taxonomy. For example, Zhao et al. (2023a) and Sato et al. (2023) address both Text Quality and Model Ownership Verification, preserving readability while ensuring ownership with detectable markers. Similarly, the work by Yang et al. (2023a) is listed under both message injection and embedding-level addition, as it involves injecting watermarks as messages within the embedding space. Similarly, research presented by Peng et al. (2023) falls under both embedding-level additions and trigger-set because it uses modified embeddings activated by specific triggers to enhance watermark detectability. These dual listings reflect the flexibility provided by these techniques, addressing overlapping goals. In contrast to prior surveys with limited focus areas, we believe maintaining separate but overlapping categories helps clarify distinct objectives and evaluation criteria across multiple dimensions, ensuring comprehensive coverage of the taxonomy.
@@
Text Completion Colossal Clean Crawled Corpus (C4) Raffel et al. (2020), Dbpedia Class Auer et al. (2007), WikiText-2 Merity et al. (2016) Kirchenbauer et al. (2023), Kuditipudi et al. (2023), Liu et al. (2023a), Munyer and Zhong (2023), Yoo et al. (2023b), Liu et al. (2023b), Fairoze et al. (2023), Ren et al. (2023), Hou et al. (2023), Qu et al. (2024), Wang et al. (2023), Liu and Bu (2024), Chen et al. (2023), Mao et al. (2024), Chang et al. (2024), Zhou et al. (2024)
Post-watermark text similarity analysis WikiText-2, Workshop on Statistical Machine Translation (WMT14) Bojar et al. (2014), Internet Movie Database (IMDb) Maas et al. (2011), AgNews Zhang et al. (2015), Dracula, Pride and Prejudice, Wuthering Heights Gerlach and Font-Clos (2020), CNN/Daily Mail Nallapati et al. (2016), Human ChatGPT Comparison Corpus (HC3) Guo et al. (2023), C4, Reuters Corpus Lewis et al. (2004), ChatGPT Abstract Nicolai Thorer Sivesind (2023), Human Abstract Nicolai Thorer Sivesind (2023) Yang et al. (2022), He et al. (2022a), He et al. (2022b), Yoo et al. (2023a), Sato et al. (2023), Zhang et al. (2023), Yang et al. (2023b), Topkara et al. (2006a)
@@
Code Generation CodeSearchNet (CSN) Husain et al. (2019), HUMANEVAL Chen et al. (2021), Mostly Basic Python Programming (MBPP), MBXP Athiwaratkun et al. (2023), DS-1000 Lai et al. (2023), APPS Hendrycks et al. (2021) Li et al. (2023), Yang et al. (2023a), Guan et al. (2024), Lee et al. (2023), Li et al. (2024), Mao et al. (2024)
Question Answering OpenGen Krishna et al. (2024), Long Form Question Answering (LFQA) Krishna et al. (2024), TruthfulQA Lin et al. (2021) Zhao et al. (2023a), Yoo et al. (2023b), Qu et al. (2024), Zhou et al. (2024), Chang et al. (2024), Chen et al. (2023)
@@ 2.3 Evaluation
A wide variety of datasets have been used to evaluate the performance of watermarking approaches, limiting our ability to extract generalized conclusions about their performance. Different benchmarks focus on selected downstream tasks to validate watermarking capabilities, and we provide a detailed breakdown of the datasets utilized in Table 3. We observe many evaluation datasets focusing on text completion and post-watermarking text similarity tasks. The downstream task descriptions are provided below.
@@
Text Completion Task
This task involves giving the LLM a portion of text from the dataset as a prompt and then asking it to complete the text. The generated completion is then compared with the human completion or the portion of the dataset not provided as the prompt.
Post-watermark text similarity analysis
In this task, given an initial text $X$, watermarking is applied to $X$ to produce a modified text $X^{\prime}$. An example could be a rule-based substitution with synonyms or spelling replacements. The comparison is then made between $X$ and $X^{\prime}$ based on distinctions in length, semantics, and other linguistic features.
Other Downstream Tasks
For these tasks, given the same initial prompt $X$, the LLM’s generated response $Y$ (before watermarking) is compared with the response $Y^{\prime}$ (after watermarking).
@@ 2.4 Adversarial attacks on watermarking techniques
Figure 4: An example of an adversary performing a de-watermarking attack on a green-red list-based watermarking technique. The original partitioning contains a higher proportion of green tokens than the partitioning after adversarial paraphrasing.
@@
    Text insertion attacks involve adding additional tokens or text segments to the original output of a watermarked LLM generation. For example, on watermarking methods with green-red lists Kirchenbauer et al. (2023); Zhao et al. (2023b); Takezawa et al. (2023), an attacker could add additional tokens from the red list, leading to the obfuscation of the watermarking method. Another variant of text insertion attacks includes copy-paste attacks Qu et al. (2024), where adversaries insert copied text from external sources into the watermarked output. This approach can dilute the effect of the watermark by embedding non-watermarked content within the watermarked text, further complicating watermark detection and attribution.
@@
    Text deletion attacks involve the removal of tokens or text segments from the original watermarked output of an LLM and modifying the rest of the tokens to fit the output. Returning to the example of green-red list methodologies, this means removing some of the green list tokens from the output and modifying the red list tokens in the output Kirchenbauer et al. (2023). These techniques often require knowledge of the vocabularies belonging to each of the two lists in green-red lists.
@@ 3 Discussion and Open Challenge
Our taxonomy-driven categorization of the research space exposes the open challenges to watermarking and outlines "good to have" criteria while developing new techniques to protect intellectual property ownership. They are as follows:
@@
One of the critical challenges in the field is the lack of comprehensive evaluation of techniques against a diverse range of de-watermarking attacks. While many researchers focus on developing robust techniques, there is often insufficient emphasis on systematic red-teaming of these methods against multiple attacking scenarios.
@@
Watermarks modify the model output distributions; techniques that are robust to de-watermarking often have greater variations in watermarked outputs compared to clean outputs, leading to a potential trade-off between de-watermarking and LLM factuality. Despite this potential trade-off, there is a lack of analysis on how watermarking techniques affect the output inaccuracies or hallucinations. After training or fine-tuning LLMs with specific watermarking techniques, there is often insufficient examination of whether these methods introduce or exacerbate inaccuracies. We advocate for factuality evaluations post-watermarking.
Enhanced Interpretability
@@
We urge the community to work on the human perception of LLMs when interacting with different safety principles. User perception of LLMs may change with differences in output distributions. Furthermore, safety practices may enable AI acceptance and adoption among the masses.
Personalized watermarking
Developing personalized watermarking approaches that cater to individual user preferences, contexts, or specific organizational policies is important. This requires adaptive techniques that can embed user- or context-specific identifiers while respecting privacy, personalization preferences, and practical constraints.
Multi-agent Watermarking
Current watermarking methods do not effectively support scenarios where multiple LLM agents sequentially or collaboratively contribute to generating text. Specifically, existing approaches fail to maintain watermark integrity and detectability when multiple distinct watermarks are stacked or embedded in succession. An unresolved challenge is developing watermarking techniques that enable each agent to reliably detect its own watermark and a collaboratively embedded watermark, even when multiple watermarks from different agents coexist within the same generated content.
@@ 4 Conclusion
In this paper, we analyze representative literature to provide a comprehensive taxonomy for digital watermarking techniques for both LLM-generated and human-written text. The taxonomy categorizes watermarking techniques using four primary categories, namely - intention of the method, data used for evaluation, watermark addition, and adversarial attacks.
We identify and cluster existing watermarking methods, highlighting key open challenges and research gaps in the field. For every watermarking method, we advocate for establishing stronger evaluation paradigms: standardized datasets, resilience to adversarial attacks, impact on model utility and output actuality, enhanced interpretability, and human perception change upon the use of these such techniques. We envision this research as a reference for policymakers, safety practitioners, and end users, facilitating the adoption of robust digital watermarking practices and promoting responsible AI use.
@@ 5 Limitations
Limitations to our work are as follows: (1) We do not include detailed insights into metrics for success rate (accuracy of detecting watermarked texts), text quality (perplexity and semantics), NLP task-specific evaluation, and robustness (detectability of watermarks after removal attacks). However, we briefly describe the two types of metrics into watermarking success rates (intrinsic quality of watermarking) and model utility/ performance 8.2. (2) Given the scope of this paper, we do not demonstrate the mathematical analysis of different watermarking techniques. We urge readers to refer to the original papers for the same (3) We do not cover all different task deployment scenarios for the watermarking techniques discussed. (4) Readers may perceive similarity between sections 2.1.1 (text quality) and 2.3 (evaluation), however, we wish to highlight the following distinction - in section 2.3 (evaluation), our focus is on the datasets that different papers use to evaluate their watermarking techniques whereas in section 2.1.1 (text quality), we look at how some papers have similar intentions or end goals of watermarking. While similar intentions often include many of the same downstream tasks as mentioned in sections 2.3 (Text Completion Task, Post-watermark text similarity analysis, Other Downstream Tasks), and 2.1.1, we advocate for the standardization of these tasks for definitive apples to apples comparison between techniques. (5) While we advocate for a standardized evaluation, we do not propose a framework for evaluating the effectiveness of watermarking techniques. (6) We focus primarily on watermarking and do not delve into the broader relationships between watermarking and adjunct fields, for example, steganography. However, additional information on these topics can be found in the appendix section 8.1.
@@ 6 Ethical Considerations
7 Acknowledgments
This work was partly supported by U.S. NSF awards #1934782 and #2114824.